# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ 4.9.3 Stellar masses and SFRs
In blue, green, and orange, we show estimates of stellar mass and SFR from JWST NIRSpec spectropscopy and NIRCam photometry for a selection of luminous galaxies from the JWST Advanced Deep Extragalactic Survey (JADES Eisenstein et al., 2023), Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2023a, ERS-1345), and (COSMOS-Web Casey et al., 2023, GO#1727). We source our data for each survey from Harikane et al. (2024), Carniani et al. (2024), and Casey et al. (2024) respectively. This selection includes the most luminous and high redshift galaxies spectroscopically confirmed at the time of writing, such as GS-z11-0, GS-z14-0, and GS-z14-1.
@@ 5 Conclusions and Future Work
In this work, we run zoom-in simulations of a MDG ($M_{\rm halo}=10^{11}~{}M_{\odot}$, $z=9$) using the cosmological hydrodynamics code RAMSES at an effective resolution $\simeq 10~{}{\rm pc}$. Our simulations capture the unique regime of high densities $\Sigma_{\rm gas}\sim 10^{4}~{}{\rm M_{\odot}/pc^{2}}$, $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ where star formation is expected to be enhanced. Our suite of simulations vary in the details of their star formation and feedback recipes, allowing us to investigate the connection between global SFE and local gas conditions. Our fiducial simulation uses a physically-motivated, turbulence-based, multi-freefall model of star formation, avoiding ad hoc extrapolation from lower redshifts.
@@
  • Star formation efficiencies: The high densities $\Sigma_{\rm gas}\sim 10^{4}~{}{\rm M_{\odot}/pc^{2}}$, $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ in star-forming regions of MDGs naturally result in efficient star formation, with a local SFE $\epsilon_{\rm ff}\sim 10-20\%$ and an integrated SFE $\epsilon_{\rm int}\sim 10-30\%$.
  • Galaxy structure: Our simulated galaxies form a thick, rotationally-supported disc. For cold gas, the disc radius is $R_{\rm gas}\sim 700~{}{\rm pc}$ and the scale height is $H/R\sim 0.5$. For stars the disc radius is $R_{\rm gas}\sim 400~{}{\rm pc}$ and scale height is $H/R\sim 0.25$. The gas scale height calculation for the disc may be artificially inflated by accretion of cold gas. The baryonic mass is organized into dense, cold, and turbulent clumps of diameter $\sim 100~{}{\rm pc}$ where most of the star formation occurs, although the size of these clumps may be an artifact of our $\sim 10~{}{\rm pc}$ resolution. These discs may have analogues in the discs of star-forming galaxies at cosmic noon, which are better studied in theory and observation.
  • Star formation history and variability The stellar mass in the MDG increases in proportion to the accretion rate onto the halo i.e. nearly exponential growth for $z\geq 9$. The SFR fluctuates on various timescales. The Fourier transform of the SFR may encode information about the typical density of star formation $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ and the typical lifetime of star-forming clouds $\tau_{\rm cl}\sim 10~{}{\rm Myr}$.
  • Stellar feedback: Our implementation of stellar feedback is crude, but our preliminary results suggest that stellar feedback in MDGs is weak due to the high density conditions. This is consistent with the general prediction of the FFB scenario, despite the fact that our current simulations don’t yet resolve the SN cooling radius and the star-forming clusters where FFB is expected to occur. Weak feedback can be partially mitigated by photoionization from massive stars, which decrease the density around SN explosions. However, even in our strongest photoionization feedback model the integrated SFE is still high $\geq 10\%$.
@@
  • Turbulence: Turbulence generally decreases the local SFE for the same gas conditions due to turbulent pressure support. The strong turbulence $\mathcal{M}_{\rm turb}\sim 30$ in MDGs prevents star-forming clouds from catastrophically collapsing. However, turbulence is not strong enough to take MDGs out of the self-regulated regime, so it has only a weak effect on global and integrated SFEs.
  • Matter-radiation anti-correlation Radiation tends to move through low-density channels in the density field, so strong turbulence decreases the effective optical depth by a factor proportional to $\mathcal{M}_{\rm turb}^{-2}$. This effect may reduce the ability for radiation pressure feedback to suppress star formation and cause SED fitting codes to overestimate stellar masses and SFRs.
@@
Our study is the first step towards understanding efficient star formation in MDGs. Our treatment of stellar feedback is crude due to the limitations of our effective resolution $\sim 10~{}{\rm pc}$ in a pure hydrodynamics simulation (i.e. no radiation). Although these factors affect the accuracy of our results, they also make the simulations easier to interpret by simplifying the physics. We will use the insights from this work to make informed choices about more computationally expensive future simulations which capture more of the physics.
The interplay between photoionization and SN feedback in our simulations suggests that early feedback is important in MDGs, even when it cannot suppress star formation directly. In future simulations, we will model early feedback in more detail using RAMSES-RT, including explicit models for photoionization, radiation pressure, and stellar winds. These models should account for the effect of strong turbulence, which can decrease the effective opacity (Sec. 4.9.2). Using RAMSES-RT will also enable a more self-consistent treatment of the extragalactic radiation field.
The high densities in our simulations mean that we do not resolve the scale of star-forming clouds, which can be roughly approximated by the Jeans radius $R_{\rm J}=c_{\rm s}/\sqrt{G\rho}$. For a cloud at the Hydrogen ionization temperature $T\sim 10^{4}~{}{\rm K}$ and a density given by the median star-forming density of our fiducial simulation $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, the Jeans radius is $R_{\rm J}\sim 20~{}{\rm pc}$, and the cloud will only be resolved by eight cells $\Delta x_{\rm min}\sim 10~{}{\rm pc}$. For a lower temperature given by the median star-forming temperature of our fiducial simulation $T\sim 300~{}{\rm K}$, the Jeans radius is $R_{\rm J}\sim 3~{}{\rm pc}$, and the cloud will not be resolved at all. The addition of 3 AMR levels would give $\Delta x_{\rm min}\sim 1~{}{\rm pc}$, resulting in fully resolved star-forming clouds. A high resolution would also solve the problem of unresolved HII regions, revealing the multi-phase structure of the ISM.
@@ Acknowledgements
Z.A. would like to acknowledge helpful conversations with David Setton, Jenny Greene, Ronan Hix, Nicholas Choustikov, and Matthew Sampson and work by Diego Solorio on the simulation initial conditions.
This material is based upon work supported by the National Science Foundation (NSF) and the U.S.-Isreal Binational Science Foundation (BSF) under Award Number 2406558 and Award Title “The Origin of the Excess of Bright Galaxies at Cosmic Dawn”.
@@
The simulations in this work were run on Stellar cluster at Princeton University. Each simulation contains approximately $1.3~{}{\rm Tb}$ of data. Access to this data can be provided upon request to the corresponding author.
@@
Appendix A Simple estimate for baryon surface density
@@
and a halo radius
$R_{\rm halo}=\left(\frac{M_{\rm halo}}{(4\pi/3)\rho_{\rm halo}}\right)^{1/3}=\left(\frac{M_{\rm halo}}{(4\pi/3)\Delta\rho_{\rm crit,0}\Omega_{\rm m,0}}\right)^{1/3}(1+z)^{-1}$ (52)
where $\rho_{\rm crit,0}=3H_{0}^{2}/(8\pi G)$ is the critical density of the Universe at redshift $z=0$.
@@
$R_{\rm gal}=\lambda R_{\rm halo}$ (53)
Based on tidal-torque theory, Dekel et al. (2013) argue that the average $\lambda$ is similar to the halo spin parameter, with typical value $\lambda\simeq 0.05$ (e.g. Bullock et al., 2001).
We further assume that the baryon mass in the galactic disc approximately equal to the total accreted baryon mass,
$M_{\rm gas}=f_{\rm b}M_{\rm halo}$ (54)
Under these assumptions, the baryon surface density $\Sigma_{\rm gas}$ in a galaxy is given by
$\Sigma_{\rm gas}=\frac{M_{\rm gas}}{\pi R_{\rm gal}^{2}}=\frac{f_{\rm b}M_{\rm halo}^{1/3}}{\pi\lambda^{2}}\left(\frac{4\pi}{3}\Delta\rho_{\rm crit,0}\Omega_{\rm m,0}\right)^{2/3}(1+z)^{2}$ (55)
We use Equation 55 to estimate the baryon surface density in a galaxy as a function of redshift and the mass of its host halo in Figure 1.
@@ Appendix C Calculation of net angular momentum
    Determine the radius $R$ such that $M_{\rm cyl}(H,R)=fM_{\rm cyl}(H,R_{\rm max})$.
@@
We set the initial disc height $H=500~{}{\rm pc}$ and the upper bound on the disc radius $R_{\rm max}=1~{}{\rm kpc}$. However, our results are not sensitive to changes in these parameters by factors of a few. We chose a mass fraction $f=85\%$ consistent with Mandelker et al. (2014). The calculation can be repeated separately for cold gas and stars. We only include gas with temperatures $T/\mu\leq 10^{4}~{}{\rm K}$ to exclude outflows and accreting gas.
@@ Appendix D Calculation of summary statistics
In Section 3.5, we used various summary statistics to compare our simulations, including the total stellar mass $M_{*}$ in the galaxy at $z=9$, the average SFR over the last $50~{}{\rm Myr}$ at $z=9$ ${\rm SFR}_{50}$, the integrated SFE $\epsilon_{\rm int}$, the median local SFE ${\rm med}(\epsilon_{\rm ff})$, and the outflow efficiency $\eta$. We compute each statistic using a box $\mathcal{B}$ of side length $5~{}{\rm kpc}$ centered on the galaxy. In our simulations, this box contains $\sim 1/3$ of the total stellar mass in the entire volume.
We compute the stellar mass by summing the star particle mass within the box at $z=9$:
$M_{*}=\sum_{\mathbf{r}_{k}\in\mathcal{B}}m_{k}$ (59)
where $\mathbf{r}_{k}$ and $m_{k}$ are the position and mass respectively of star particle $k$.
The average SFR over the last $50~{}{\rm Myr}$ at $z=9$ (denoted here as ${\rm SFR}_{50}$) is given by
[…]
where $t_{9}$ is the proper time at $z=9$.
We compute the SFR as a function of time by binning the star particle mass within the box by time of birth:
$\widetilde{\dot{M}}_{*}(t)=\frac{1}{\Delta t}\sum_{\begin{subarray}{c}\mathbf{x}_{k}\in\mathcal{B}\\ t_{{\rm birth},k}\in[t,t+\Delta t]\end{subarray}}m_{k}$ (61)
where $t_{{\rm birth},k}$ is the time of birth for star particle $k$. We use 5500 time bins of width $\Delta t=100~{}{\rm kyr}$ in the range $t\in[0,550]~{}{\rm Myr}$.
Equation 61 underestimates the true SFR because the star particle mass decreases over its lifetime due to SN mass loss. After adjusting for SN mass loss, the SFR is
$\dot{M}_{*}(t)=\frac{\widetilde{\dot{M}}_{*}(t)}{f(t_{z=9}-t)}$ (62)
@@
The local SFE is a function of the local gas conditions by the MFF model (Eq. 22). We compute the median local SFE from the distribution of local SFEs in star birth events, weighted by stellar mass. The integrated SFE is defined by
@@
We compute the stellar mass by integrating the SFR over time, accounting for the mass loss due to SNe:
@@
where $f(\Delta t)$ is defined by Equation 63.
The outflow efficiency $\eta$ is defined by the ratio of mass outflow to SFR $\eta=\dot{M}_{\rm out}/\dot{M}_{*}$. The outflow rate is computed as the mass flux of outflowing gas in a spherical shell $\partial V$ of radius $R_{\rm out}=2~{}{\rm kpc}$ and thickness $\Delta R_{\rm out}=0.1R_{\rm out}=200~{}{\rm pc}$:
[…]
where $H$ is the Heaviside step function, $v_{r}$ is the radial velocity of the gas in a coordinate system where the origin is at the center of the galaxy and $\Delta V_{\rm out}$ is the volume of the spherical shell, given by
@@
For each simulation, we compute the outflow rate at $z=9$ and divide by ${\rm SFR}_{50}$ to get the outflow efficiency. For a given simulation, the value of $\eta$ fluctuates for different choices of $R_{\rm out}$, but the ratio of outflow efficiencies between simulations is robust for values $R_{\rm out}\in[1,4]~{}{\rm kpc}$.

### v2

@@ 4.9.2 Stellar masses and SFRs
In blue, green, and orange, we show estimates of stellar mass and SFR from JWST NIRSpec spectropscopy and NIRCam photometry for a selection of luminous galaxies from the JWST Advanced Deep Extragalactic Survey (JADES Eisenstein et al., 2023), Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2024, ERS-1345), and (COSMOS-Web Casey et al., 2023, GO#1727). We source our data for each survey from Harikane et al. (2024), Carniani et al. (2024), and Casey et al. (2024) respectively. This selection includes the most luminous and high redshift galaxies spectroscopically confirmed at the time of writing, such as GS-z11-0, GS-z14-0, and GS-z14-1.
@@ 5 Conclusions and Future Work
In this work, we run zoom-in simulations of a MDG ($M_{\rm halo}=10^{11}~{}M_{\odot}$, $z=9$) using the cosmological hydrodynamics code RAMSES at an effective resolution $\simeq 10~{}{\rm pc}$. Our simulations capture a unique regime where star formation is expected to be enhanced by high gas densities. Our suite of simulations vary in the details of their star formation and feedback recipes, allowing us to investigate the connection between global SFE and local gas conditions. Our fiducial simulation uses a physically-motivated, turbulence-based, multi-freefall model of star formation, avoiding ad hoc extrapolation from lower redshifts.
@@
  • Star formation efficiencies: The high gas densities $\sim 10^{4}~{}{\rm M_{\odot}/pc^{2}}$, $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ in star-forming regions of MDGs naturally result in efficient star formation, with a local SFE $\epsilon_{\rm ff}\sim 10-20\%$ and an integrated SFE $\epsilon_{\rm int}\sim 10-30\%$.
  • Galaxy structure: Our simulated galaxies form a thick disc with support from both rotation and turbulence. For cold gas, the disc radius is $R_{\rm gas}\sim 700~{}{\rm pc}$ and the scale height is $H/R\sim 0.5$. For stars the disc radius is $R_{\rm gas}\sim 400~{}{\rm pc}$ and scale height is $H/R\sim 0.25$. The baryonic mass is organized into dense, cold, and turbulent clumps of diameter $\sim 100~{}{\rm pc}$ where most of the star formation occurs, although the size of these clumps may be an artifact of our $\simeq 10~{}{\rm pc}$ effective resolution. These discs may have analogues in the discs of star-forming galaxies at Cosmic Noon.
  • Star formation history and variability: The stellar mass in the MDG increases in proportion to the accretion rate onto the halo i.e. nearly exponential growth for $z\geq 9$. The SFR fluctuates on various timescales, producing an intrinsic UV variability $\sigma_{\rm UV}\simeq 0.83$ which is reduced to $\sigma_{\rm UV}\simeq 0.35$ when measured with a wide-band filter. This is insufficient to explain the discrepancies between models and data from bursty star formation alone. The Fourier transform of the SFR encodes information about the typical density of star formation $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ and the typical lifetime of star-forming clouds $\sim 10~{}{\rm Myr}$.
  • Stellar feedback: Our implementation of stellar feedback is crude, but our preliminary results suggest that stellar feedback in MDGs is weak due to the high density conditions. This is consistent with the general prediction of the FFB scenario, despite the fact that our current simulations don’t resolve the SN cooling radius or the star-forming clusters where FFBs would occur. Weak feedback can be partially mitigated by photoionization from massive stars, which decreases the density around SN explosions. However, even in our strongest photoionization feedback model the integrated SFE is still high $\geq 10\%$.
@@
  • Turbulence: Turbulence generally decreases the local SFE for the same gas conditions due to turbulent pressure support. Turbulence in MDGs is strong $\sim 50~{}{\rm km/s}$ due to intense accretion, but not strong enough to take MDGs out of the self-regulated regime, so it has only a weak effect on global and integrated SFEs.
@@
Our study is the first step towards understanding efficient star formation in MDGs. Our treatment of stellar feedback is crude due to the limitations of our effective resolution $\simeq 10~{}{\rm pc}$ in a pure hydrodynamics simulation. Although these factors affect the accuracy of our results, they also make the simulations easier to interpret by simplifying the physics. We will use the insights from this work to make informed choices about more computationally expensive future simulations which capture more of the physics.
The interplay between photoionization and SN feedback in our simulations suggests that early feedback is important in MDGs, even when it cannot suppress star formation directly. In future simulations, we will model early feedback in more detail using RAMSES-RT, including explicit models for photoionization, radiation pressure, and stellar winds. These models should account for the effect of strong turbulence, which can decrease the effective opacity (Sec. LABEL:sec:turb_deatt). Using RAMSES-RT will also enable a more self-consistent treatment of the extragalactic radiation field.
The high densities in our simulations mean that we do not resolve the scale of star-forming clouds, which can be roughly approximated by the Jeans radius $R_{\rm J}=c_{\rm s}/\sqrt{G\rho}$. For a cloud at the Hydrogen ionization temperature $\sim 10^{4}~{}{\rm K}$ and a density given by the median star-forming density of our fiducial simulation $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, the Jeans radius is $\sim 20~{}{\rm pc}$, and the cloud will only be resolved by eight cells. For a lower temperature given by the median star-forming temperature of our fiducial simulation $\sim 300~{}{\rm K}$, the Jeans radius is $\sim 3~{}{\rm pc}$, and the cloud will not be resolved at all. The addition of 3 AMR levels would give an effective resolution $\simeq 1~{}{\rm pc}$, resulting in fully resolved star-forming clouds. A high resolution would also solve the problem of unresolved HII regions, revealing the multi-phase structure of the ISM.
@@ Acknowledgements
Z.A. would like to acknowledge helpful conversations with David Setton, Jenny Greene, Eve Ostriker, Ronan Hix, Nicholas Choustikov, Matthew Sampson, Diederek Kruijssen, Mike Boylan-Kolchin, Tibor Dome, Robert Feldman, and work by Diego Solorio on the simulation initial conditions. We are grateful to our referee Daniel Ceverino, whose time and attention significantly improved the quality of this manuscript.
This material is based upon work supported by the National Science Foundation (NSF) and the U.S.-Israel Binational Science Foundation (BSF) under Award Number 2406558 and Award Title “The Origin of the Excess of Bright Galaxies at Cosmic Dawn”.
@@
The simulations in this work were run on Stellar cluster at Princeton University. Each simulation contains approximately $1.3~{}{\rm Tb}$ of data. Due to data storage limitations, we have only maintained data from our fiducial simulation, which can be provided upon reasonable request to the corresponding author. Likewise, the exact RAMSES patch and initial conditions used for our simulations can be provided upon reasonable request to the corresponding author.
@@
Appendix A Simple estimate of baryon surface density
@@
and a halo virial radius
$\begin{split}R_{\rm vir}=\ &\left(\frac{M_{\rm vir}}{(4\pi/3)\rho_{\rm halo}}\right)^{1/3}=\left(\frac{M_{\rm vir}}{(4\pi/3)\Delta\rho_{\rm crit,0}\Omega_{\rm m,0}}\right)^{1/3}(1+z)^{-1}\\ \simeq\ &14M_{11}^{1/3}(1+z)_{10}^{-1}~{}{\rm kpc}\end{split}$ (54)
where $M_{11}=M_{\rm vir}/10^{11}~{}M_{\odot}$, $(1+z)_{10}=(1+z)/10$, and $\rho_{\rm crit,0}=3H_{0}^{2}/(8\pi G)$ is the critical density of the Universe at redshift $z=0$. The corresponding virial temperature is
$\frac{T_{\rm vir}}{\mu}=\frac{m_{\rm p}}{k_{\rm B}}\frac{GM_{\rm vir}}{2R_{\rm vir}}\simeq 3.6\times 10^{6}M_{11}^{2/3}(1+z)_{10}~{}{\rm K}$ (55)
@@
$R_{\rm gal}=\lambda R_{\rm vir}\simeq 720\lambda_{0.05}M_{11}^{1/3}(1+z)_{10}^{-1}~{}{\rm pc}$ (56)
where $\lambda_{0.05}=\lambda/0.05$. Based on tidal-torque theory, Dekel et al. (2013) argue that the average $\lambda$ is similar to the halo spin parameter, with typical value $\lambda\simeq 0.05$ (e.g. Bullock et al., 2001). We further assume that the baryon mass in the galactic disc is approximately equal to the total accreted baryon mass,
$M_{\rm b}=f_{\rm b}M_{\rm vir},\quad f_{\rm b}=\Omega_{\rm b,0}/\Omega_{\rm m,0}$ (57)
Under these assumptions, the typical baryon surface density in the galaxy is
$\begin{split}\Sigma_{\rm b}=\ &\frac{M_{\rm b}}{\pi R_{\rm gal}^{2}}=\frac{f_{\rm b}M_{\rm vir}^{1/3}}{\pi\lambda^{2}}\left(\frac{4\pi}{3}\Delta\rho_{\rm crit,0}\Omega_{\rm m,0}\right)^{2/3}(1+z)^{2}\\ \simeq\ &9.6\times 10^{3}\lambda_{0.05}^{-2}M_{11}^{1/3}(1+z)_{10}^{2}~{}M_{\odot}/{\rm pc^{2}}\end{split}$ (58)
As a corollary, we see that $M_{\rm vir}\propto(1+z)^{6}$, which is similar to a result from Boylan-Kolchin (2024). We use Equation 58 to estimate the baryon surface density in a galaxy as a function of redshift and the mass of its host halo in Figure 1. The gas and baryon surface densities are related by $\Sigma_{\rm gas}=(1-\epsilon_{\rm int})\Sigma_{\rm b}$, where $\epsilon_{\rm int}=M_{*}/M_{\rm b}$ is the integrated SFE. Given the disk scale height $(H/R)_{\rm gal}$, the typical number density is
$\begin{split}n_{\rm H}=\ &\frac{M_{\rm gas}/m_{\rm H}}{\pi R_{\rm gal}^{3}(H/R)_{\rm gal}}=\frac{(1-\epsilon_{\rm int})f_{\rm b}}{\pi\lambda^{3}(H/R)_{\rm gal}}\frac{4\pi}{3}\Delta\rho_{\rm crit,0}\Omega_{\rm m,0}(1+z)^{3}\\ \simeq\ &540~{}(1-\epsilon_{\rm int})\lambda_{0.05}^{-3}(H/R)(1+z)_{10}^{3}~{}{\rm cm^{-3}}\end{split}$ (59)
@@ Appendix C Calculation of net angular momentum
    Determine the radius $R$ such that $M_{\rm cyl}(H,R)=fM_{\rm cyl}(H,R_{\rm max})$;
@@
We set the initial disc height $H=500~{}{\rm pc}$ and the upper bound on the disc radius $R_{\rm max}=1~{}{\rm kpc}$. However, our results are not sensitive to changes in these parameters by factors of a few. We chose a mass fraction $f=85\%$ consistent with Mandelker et al. (2014). The calculation can be repeated separately for cold gas and stars. We only include gas with temperatures $\leq 10^{4}~{}{\rm K}$ to exclude outflows and accreting gas.
@@ Appendix D Calculation of summary statistics
We use various summary statistics to compare our simulations in Section 3.5. In this appendix, we explain the calculation details for each statistic. For these calculations, we restrict the domain to a box $\mathcal{B}$ of side length $5~{}{\rm kpc}$ centered on the galaxy to avoid contributions from other galaxies in the simulation box. At $z=9$, $\mathcal{B}$ contains $\sim 1/3$ of the stellar mass in the simulation box.
Our simulations record the time of birth for each star particle. To compute the SFR, we start by binning the star particles in $\mathcal{B}$ at $z=9$ by time of birth and dividing by the bin width:
$\widetilde{\dot{M}}_{*}(t)=\frac{1}{\Delta t}\sum_{\begin{subarray}{c}\mathbf{r}_{k}(t_{9})\in\mathcal{B}\\ t_{{\rm b},k}\in[t,t+\Delta t]\end{subarray}}m_{k}(t_{9})$ (63)
where $\mathbf{r}_{k}$ is the position of star particle $k$, $t_{{\rm b},k}$ is the time of birth of star particle $k$, $\Delta t$ is the bin width, and $t_{9}$ is the time at $z=9$. We use 5500 time bins of width $100~{}{\rm kyr}$ in the range $[0,550]~{}{\rm Myr}$.
Equation 63 underestimates the SFR because the star particle mass decreases over its lifetime due to SN mass loss. This is accounted for by
$\dot{M}_{*}(t)=\frac{\widetilde{\dot{M}}_{*}(t)}{f(t_{9}-t)}$ (64)
@@
The SFR averaged over a time interval $\Delta t$ is
[…]
We compute the stellar mass by integrating the SFR over time, accounting for SN mass loss:
[…]
Equivalently, we can compute the stellar mass by summing the masses of the star particles in $\mathcal{B}$:
$M_{*}(t)=\sum_{\mathbf{r}_{k}(t)\in\mathcal{B}}m_{k}(t)$ (68)
In the fiducial simulation, the two methods always agree to within 1%. However, we prefer Equation 67 because it can be used to compute the stellar mass at arbitrary times rather than only times associated with data dumps. We compute the gas mass by summing the mass $\rho\Delta V$ in each grid cell, similar to Equation 68.
To compute the virial mass, we start by computing the enclosed mass $M(<r)$ as a function of radius, including contributions from gas, stars, and dark matter. The radial coordinate is computed with respect to the center of the halo. The virial radius is the radius where $M(<R_{\rm vir})=(4/3)\pi R_{\rm vir}^{3}\rho_{\rm halo}$, where $\rho_{\rm halo}$ is given by Equation 53. The virial mass is $M_{\rm vir}=M(<R_{\rm vir})$.
The local SFE is a function of the local gas conditions by the MFF model (Eq. 22). We compute the median local SFE from the distribution of local SFEs in star birth events, weighted by stellar mass. The integrated SFE is
@@
The outflow efficiency is given by $\eta=\dot{M}_{\rm out}/{\rm SFR}_{\rm 50}$, where $\dot{M}_{\rm out}$ is the mass outflow rate and ${\rm SFR}_{\rm 50}$ is the SFR averaged over the last $50~{}{\rm Myr}$. We use ${\rm SFR}_{50}$ rather than $\dot{M}_{*}$ because the raw SFR is highly stochastic, making the mass outflow rate better-correlated with ${\rm SFR}_{\rm 50}$. We compute the mass outflow rate using the mass flux of gas with positive radial velocities in a spherical shell $\partial V$ of radius $R_{\rm out}=2~{}{\rm kpc}$ and thickness $\Delta R_{\rm out}=0.1R_{\rm out}=200~{}{\rm pc}$:
@@
where $H$ is the Heaviside step function, $v_{r}$ is the radial velocity of the gas relative to the center of the galaxy, and $\Delta V_{\rm out}$ is the volume of the spherical shell, given by
@@
For a given simulation, the value of $\eta$ varies for different choices of $R_{\rm out}$, but the ratio of outflow efficiencies between simulations is robust for values from $1$ to $4~{}{\rm kpc}$.